• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 09:09:05 on July23,2003

    Entry number 106310

    keyword=Target cell replaced.

    Vince and I installed a new target ladder with a new He3 cell (Priapus). Also installed a new reference cell, new 12C targets (short ones at -20cm (upstream) and +10 cm
    (downstream), long one at the center). The BeO target is also installed so that when beam come back we can center beam on the BeO target. Also the left arm 12C target
    is removed to get ready for optics run. Thin sieve slits inserted. Target collimators removed.

    Cooling jets turned back on. Air flow turned back on, started oven heating.

    We removed as much as we could the ice inside the sieve box near the entrance of the Septum.